OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 3000 fewer panels than Plant B did. Four percent of the panels from Plant A and 3% of the panels from Plant B were defective. How many panels did Plant A produce, if the two plants together produced 1070 defective panels?

OpenStudy (mathstudent55):

The first step is to choose two variables to represent the numbers of panels made in each plant.

OpenStudy (anonymous):

A= b-3000

OpenStudy (mathstudent55):

Good. You already have the two variables and the first equation.

OpenStudy (mathstudent55):

Now we need a second equation.

OpenStudy (mathstudent55):

How do you write 4% of A?

OpenStudy (anonymous):

A.04

OpenStudy (mathstudent55):

0.04A

OpenStudy (anonymous):

a*.04

OpenStudy (anonymous):

oh okay

OpenStudy (mathstudent55):

Now do the same for 3% of B

OpenStudy (anonymous):

B.03

OpenStudy (mathstudent55):

0.03B is better.

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

The numbers of defective panels are 0.04A and 0.03B, right?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Now add them together and set the sum equal to 1070. That's the second equation.

OpenStudy (anonymous):

.07=1070

OpenStudy (mathstudent55):

No. 0.04A and 0.03B are unlike terms. They can't be combined together. Their sum must be written as 0.04A + 0.03B

OpenStudy (mathstudent55):

The second equation is: 0.04A + 0.03B = 1070

OpenStudy (anonymous):

correct

OpenStudy (mathstudent55):

Now you have two equations: A = B - 3000 0.04A + 0.03B = 1070 Now you need to solve the system of equations.

OpenStudy (mathstudent55):

Since the first equation is already solved for A, we can use substitution easily.

OpenStudy (mathstudent55):

Replace A of the second equation with B - 3000. Then you will have an equation with only B, and you can solve it for B.

OpenStudy (anonymous):

.04A(b-3000)+.03=1070

OpenStudy (anonymous):

no

OpenStudy (anonymous):

.04(b-3000)+.03=1070

OpenStudy (mathstudent55):

There is still a B with 0.03

OpenStudy (anonymous):

yes so 7b-120=1070?

OpenStudy (anonymous):

b=170

OpenStudy (anonymous):

.07

OpenStudy (mathstudent55):

\(\color{red}{A = B - 3000}\) \(0.04\color{red}{A} + 0.03B = 1070\) \(0.04\color{red}{(B - 3000)} + 0.03B = 1070\)

OpenStudy (anonymous):

17000

OpenStudy (anonymous):

b=17000?

OpenStudy (mathstudent55):

0.04B - 120 + 0.03B = 1070 0.07B - 120 = 1070 0.07B = 1190 B = 17,000

OpenStudy (anonymous):

a= 14000?

OpenStudy (mathstudent55):

Correct. A = B - 3000 A = 17,000 - 3000 A = 14,000

OpenStudy (mathstudent55):

You are correct.

OpenStudy (anonymous):

YAY

OpenStudy (mathstudent55):

Answer: Plant A produced 14,000 panels.

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

You're welcome.